What volume of each solution contains 0.15 mol of KCl?

0.220 M KCl

Express your answer using two significant figures.

1.7 M KCl

Express your answer using two significant figures.

0.835 M KCl

Express your answer using two significant figures.
Answer #1

Given in the question that number of moles of KCl is 0.15 mol.

We need to find volume o each solution for the given molarities of KCL

The number of moles of solute per liter o solution is called Molarity. It is denoted by M.

M = number of moles / liters solution

(1) When molarity is 0.220 M KCl

Number of moles of solute (KCl) = 0.15 mol

Molarity of KCl = 0.220 M

We need to find volume.

So substitute the values in molarity formula, we get

0.220 = 0.15 / liters solution

So liters solution = 0.15/0.220

liters solution = 0.681818 L = 0.69 L

For 0.220 M KCl, volume of solution containing 0.15 mol of KCl is 0.69 L

(2) When molarity is 1.7 M KCl

Number of moles of solute (KCl) = 0.15 mol

Molarity of KCl = 1.7 M

We need to find volume.

So substitute the values in molarity formula, we get

1.7 = 0.15 / liters solution

So liters solution = 0.15/1.7

liters solution = 0.0823 L =0.08 L

For 1.7 M KCl, volume of solution containing 0.15 mol of KCl is 0.08 L

(3) When molarity is 0.835 M KCl

Number of moles of solute (KCl) = 0.15 mol

Molarity of KCl = 0.835 M

We need to find volume.

So substitute the values in molarity formula, we get

0.835 = 0.15 / liters solution

So liters solution = 0.15/0.835

liters solution = 0.1796 L = 0.17 L

For 0.835 M KCl, volume of solution containing 0.15 mol of KCl is 0.17 L